OpenStudy (anonymous):

why is loga a = 1 false if a is a real number?

ganeshie8 (ganeshie8):

who says it is false ?

OpenStudy (freckles):

there are certain restrictions to a

ganeshie8 (ganeshie8):

oh right, it need not be true always...

OpenStudy (anonymous):

true or false: \[\log_{a} a=1\] if \[a\]

OpenStudy (anonymous):

*if a is a real number

OpenStudy (anonymous):

and the answer is false...

ganeshie8 (ganeshie8):

\[\log_a a = 1 \iff a = a^1\] see any real numbers, \(a\), that might break above equality ?

OpenStudy (poopsiedoodle):

o well dang I was about to do a big ol' explanation there because I thought you made a typo but now I have no idea :<

jimthompson5910 (jim_thompson5910):

The domain of log(x) is x > 0

OpenStudy (freckles):

\[\log_a(a)=\frac{\ln(a)}{\ln(a)}\]

OpenStudy (freckles):

consider the domain of that expression

ganeshie8 (ganeshie8):

it seems \(a=0\) may be an issue

OpenStudy (anonymous):

so it doesn't work whenever a is equal to 0...

OpenStudy (freckles):

\[\frac{\ln(a)}{\ln(a)} \\ \text{ we want } \ln(a) \neq 0 \text{ and we want } a>0\]

OpenStudy (poopsiedoodle):

why wouldn't it work if a is 0 tho? because 0^0 still = 1

jimthompson5910 (jim_thompson5910):

a = 1 doesn't work either (plug a = 1 into the equation @freckles wrote)

jimthompson5910 (jim_thompson5910):

0^x = 0 where x is nonzero x^0 = 1 where x is nonzero 0^0 is one of many indeterminate forms

OpenStudy (freckles):

the intersections of the set of number a such that ln(a) not equal to 0 and a>0 is...

ganeshie8 (ganeshie8):

Yes, \(\log_a a=1\) holds for all real \(a\) except \(\{0,1\}\)

OpenStudy (freckles):

I don't think it holds for the negative numbers

ganeshie8 (ganeshie8):

why not

OpenStudy (freckles):

because the domain of ln(x) is x>0

OpenStudy (freckles):

i guess it depends on what class this is

OpenStudy (anonymous):

I still don't quite get it... because \[1=1^{1}\] and \[0=0^{1}\]

ganeshie8 (ganeshie8):

that is a good question, i have the same question too

OpenStudy (anonymous):

this is calculus 1 class

ganeshie8 (ganeshie8):

\(\log_0 0\) is not well defined because, it could be any thing. you may also say : \[\log_0 0 = 2 \]

ganeshie8 (ganeshie8):

or \(\log_0 0 = 3 \) or anything you wish because \(0=0^n\) for all finite nonzero \(n\)

ganeshie8 (ganeshie8):

i think, thats the reason we let \(\log_0 0\) be undefined

OpenStudy (anonymous):

ok, thank you everyone! I think I got it...

jimthompson5910 (jim_thompson5910):

Using @ganeshie8 's reasoning, \(\Large \log_{1}(1)\) is also poorly defined \(\Large \log_{1}(1) = x\) --> \(1^x = 1\) --> \(1 = 1\) where x is any real number

ganeshie8 (ganeshie8):

Exactly! I think those are the only two bad real numbers that `log` can't deal with